OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 0 and a vertical asymptote at, x=2 and x=1

OpenStudy (tkhunny):

You should be able to do these in your sleep with so few requirements. There are infintely many possibilities. \(y = \dfrac{1}{(x-2)(x-1)}\) - Done. \(y = \dfrac{1}{(x-2)(x-1)^2}\) - Done. \(y = \dfrac{4}{(x-2)(x-1)}\) - Done.